how do i solve this
find an equation of an ellipse with vertices (-3,0) and (3,0) and e= 7/10

Answer by lwsshak3(11628)

how do i solve this
find an equation of an ellipse with vertices (-3,0) and (3,0) and e= 7/10
**
Given ellipse has a horizontal major axis with equation of the standard form:
(x-h)^2/a^2+(y-k)^2/b^2=1, a>b, with (h,k) being the (x,y) coordinates of the center.
..
Solving for a, b, and coordinates of the center:
center: (0,0)
length of major axis=3-(-3)=6=2a
a=3
a^2=9
..
e=c/a=7/10=c/3
c=21/10=2.1
c^2=4.41
..
c^2=a^2-b^2
b^2=a^2-c^2=9-4.41=4.59
..
Equation of given ellipse:x^2/9+y^2/4.59=1